# What if Othello-Playing Language Models Could See?

### v1

@@ Appendix I Ablation Study
According to Table 3, both the image encoder choice and future token masking significantly affect model performance. Replacing the fine-tuned ResNet with simpler alternatives such as Pooling or Area projections results in moderate drops in legal move accuracy (from 94.03% to 92.43% and 91.80%, respectively). Using ResNet without fine-tuning also leads to lower performance (92.04%), highlighting the benefit of domain-specific adaptation. Notably, removing future token masking causes a substantial degradation (62.03%), demonstrating its critical role in aligning the training objective with the causal structure of the game.
Method   legal acc
@@
Table 3: Ablation results for VisOthello, using different image encoders (Pooling, Area, ResNet without fine-tuning) and without future token masking (W/O FTM). Results are based on the best validation checkpoints.
Appendix J Feature Alignment
@@
We treat representations from different models (e.g., Othello-GPT and VisOthello) corresponding to the same game sequence as paired training data. For example, given the Othello move sequence input “F5 F6 E6 F4 C3 D7”, the pairwise training input $H^{\prime}_{1}$ and $H^{\prime}_{2}$ correspond to the representations extracted from Othello-GPT and VisOthello models, respectively, for this exact sequence and the associated images (when applicable). The mapping matrix $W$ is learned and optimized with iterative Procrustes alignment (Gower & Dijksterhuis, 2004), which alternates between solving for the optimal orthogonal transformation and refining the mapping. This process minimizes the distance between the transformed source representations and the target representations, resulting in better alignment across the two vector spaces.
@@
We also adopt the unsupervised training approach (Conneau et al., 2018; Lample et al., 2017) with the absence of paired data or predefined anchors to learn the alignment. Given a set of game features $H^{\prime}$ from both the source and target space, the process begins with adversarial training, where a discriminator is trained to distinguish whether the feature comes from the source or target representation space. Simultaneously, the mapping matrix $W$ is optimized to make this distinction harder, effectively aligning the distributions. Once an initial mapping is obtained, we apply iterative Procrustes refinement, similar to the supervised setting, to improve the alignment. Alignment quality is evaluated and improved using the average cosine similarity between mapped source and target features on the test set.
J.3 Alignment Training Setups
@@
To construct the alignment training set, we randomly sample one subsequence from each complete game, resulting in 3,849 input sequences, each paired with the corresponding board state images. We then divide the data into training and testing sets with an 80%/20% split, resulting in 3,079 and 770 instances. We adopt cosine similarity to measure the alignment quality between representations from different models. After projecting the representations into a shared space, we compute the average pairwise cosine similarity between aligned feature vectors. A higher similarity score indicates better alignment, suggesting that the models, despite being trained on different modalities, capture similar underlying patterns. We train the alignment model using a single NVIDIA A100 GPU. All hyperparameters follow the default settings provided by the original MUSE implementation, with no additional tuning.
J.4 Mapping Result
@@
Table 4 and 5 demonstrate the mapping results under supervised and unsupervised training. We find that the alignment similarity generally improves as the size of the training data increases. This trend suggests that with more data, the models learn richer and shared representations that are easier to align across modalities. Also, despite the difference in training strategy (i.e., autoregressive training and mask language modeling), Othello-GPT and BERT exhibit strong alignment, reflected in their high similarity scores. Surprisingly, Othello-GPT exhibits a strong alignment score with VisOthello, indicating that despite differences in architecture and training modalities, the two models learn remarkably similar representations. This suggests that the underlying patterns essential for Othello gameplay are captured consistently across both language-based and multi-modal models. Such alignment highlights the potential for cross-modal knowledge transfer and opens avenues for further exploration of unified representations in complex tasks.

### v2

@@ Results.
As shown in Table 4, both the choice of image encoder and the use of FTM substantially affect performance. Replacing the fine-tuned ResNet with simple Pooling or Area projections reduces legal move accuracy from 94.03% to 92.43% and 91.80%, respectively, while using an unfine-tuned ResNet yields 92.04%. These results highlight the benefit of domain-specific adaptation for the visual encoder. More strikingly, removing FTM causes performance to collapse to 62.03%, underscoring its critical role in aligning the training objective with the causal structure of the game.
Method   Legal Move Accuracy
@@
Table 4: Ablation results for VisOthello, using different image encoders (Pooling, Area, ResNet without fine-tuning) and without future token masking (W/O FTM). All results are reported from the best validation checkpoints with training seed 42.
4.4 Probing Internal Representations
Probe training.
The linear probe is trained on 5,000 samples from the training set and evaluated on the same test set used in previous experiments. Due to architectural differences, we restrict this analysis to VisOthello and the BERT baseline. We evaluate both models after training on 0, 5k, and 20k examples, while keeping the probing setup fixed across all conditions.
Results.
Figure 2 shows F1 scores from linear probes trained to predict tile-level board states from selected layers of BERT and VisOthello. When models are randomly initialized (0 examples), VisOthello already encodes more board-relevant structure than BERT, achieving substantially higher probe performance in early layers. This may be attributed to the use of a ResNet encoder pretrained on Othello board images, which already encodes useful spatial structure. As training dataset increases, both models improve, but VisOthello consistently achieves higher scores—especially in deeper layers. After training on 20k examples, VisOthello reaches 77.55 F1 at Layer 18, compared to 62.28 for BERT. This suggests that VisOthello learns more accurate internal representations of the board state, benefiting from both multi-modal input and architectural modifications.
5 Semantically Irrelevant Perturbation
To evaluate model robustness and generalization, we test performance under semantically irrelevant perturbations—input transformations that alter surface form without changing game state.
5.1 Board Rotation
We focus on board rotation as a concrete instance, which each test board is rotated 180 degrees. As illustrated in Figure 3, this transformation corresponds to a spatial inversion: for image-based models, this involves rotating the game board in the input image; for language-based models, it requires flipping the row and column indices of the move representation (e.g., D3 becomes E6, resulting in a different move token ID).
We apply the rotation only at test time, evaluating models that were trained on the original (unrotated) training data. All models are assessed on their ability to predict both the next legal move and the exact next move, as described in Section 4. This setup allows us to examine whether models rely on absolute visual or positional cues, or whether they have learned more abstract, generalizable representations of the board state.
As shown in Figure 4, BERT remains relatively robust under board rotation, with accuracy of 90–93% across all settings. This stability is expected: since BERT operates on symbolic move sequences, board rotation can be handled through a deterministic remapping of move tokens (e.g., D3 becomes E6). In contrast, ResNet-18 suffers a substantial drop in accuracy under rotation, falling to 28–35% depending on training size and pretraining. This suggests the model fails to learn rotation-invariant representations and relies heavily on absolute spatial patterns. Lacking access to move history or turn information, ResNet depends on ambiguous visual cues that can be easily disrupted—highlighting a key limitation of purely visual models in game playing tasks like Othello.
(a) Legal move acc. with 1K samples
(b) Legal move acc. with 10K samples
(c) Legal move acc. with 20K samples
(d) Exact match acc. with 1K samples
(e) Exact match acc. with 10K samples
(f) Exact match acc. with 20K samples
Figure 4: Comparison of models’ performance with and without board rotation across different training dataset sizes. The results demonstrate that multi-modal models maintain better performance under rotation compared to purely visual models. P-P indicates the model is pretrained, while S-S indicates it is trained from scratch.
VisOthello, which combines ResNet’s image features with BERT’s move sequence encoding, maintains high accuracy after rotation (91–93%). Compared with ResNet, VisOthello receives explicit sequence information, including the player turn and previous moves, which helps disambiguate the rotated board. The language modality further guides the interpretation of visual features, enabling the model to maintain stable predictions under spatial transformations. This result illustrates the strength of multi-modal grounding: by aligning perceptual input with symbolic context, VisOthello overcomes the spatial brittleness seen in purely visual models.
6 Feature Alignment
@@ 6.2 Alignment Training
We treat representations from different models (e.g., Othello-GPT and VisOthello) corresponding to the same game sequence as paired training data. For example, given the Othello move sequence input “F5 F6 E6 F4 C3 D7”, the pairwise training input $H^{\prime}_{1}$ and $H^{\prime}_{2}$ correspond to the representations extracted from Othello-GPT and VisOthello models, respectively, for this exact sequence and the associated images (when applicable). The mapping matrix $W$ is learned and optimized with iterative Procrustes alignment (Gower and Dijksterhuis, 2004), which alternates between solving for the optimal orthogonal transformation and refining the mapping. This process minimizes the distance between the transformed source representations and the target representations, resulting in better alignment across the two vector spaces.
@@
We also adopt the unsupervised training approach Conneau et al. (2018); Lample et al. (2017) with the absence of paired data or predefined anchors to learn the alignment. Given a set of game features $H^{\prime}$ from both the source and target space, the process begins with adversarial training, where a discriminator is trained to distinguish whether the feature comes from the source or target representation space. Simultaneously, the mapping matrix $W$ is optimized to make this distinction harder, effectively aligning the distributions. Once an initial mapping is obtained, we apply iterative Procrustes refinement, similar to the supervised setting, to improve the alignment. Alignment quality is evaluated and improved using the average cosine similarity between mapped source and target features on the test set.
6.3 Alignment Training Setups
@@
To construct the alignment training set, we randomly sample one subsequence from each complete game, resulting in 3,849 input sequences, each paired with the corresponding board state images. We then divide the data into training and testing sets with an 80%/20% split, resulting in 3,079 and 770 instances. We adopt cosine similarity to measure the alignment quality between representations from different models. After projecting the representations into a shared space, we compute the average pairwise cosine similarity between aligned feature vectors. A higher similarity score indicates better alignment, suggesting that the models, despite being trained on different modalities, capture similar underlying patterns. We train the alignment model using a single NVIDIA A100 GPU. All hyperparameters follow the default settings provided by the original MUSE implementation, with no additional tuning. All models in this experiment are trained with a fixed random seed 42.
6.4 Mapping Result
@@
Table 5 and 6 demonstrate the mapping results under supervised and unsupervised training. We find that the alignment similarity generally improves as the size of the training data increases. This trend suggests that with more data, the models learn richer and shared representations that are easier to align across modalities. Also, despite the difference in training strategy (i.e., autoregressive training and mask language modeling), Othello-GPT and BERT exhibit strong alignment, reflected in their high similarity scores. Surprisingly, Othello-GPT exhibits a strong alignment score with VisOthello, indicating that despite differences in architecture and training modalities, the two models learn remarkably similar representations. This suggests that the underlying patterns essential for Othello gameplay are captured consistently across both language-based and multi-modal models. Such alignment highlights the potential for cross-modal knowledge transfer and opens avenues for further exploration of unified representations in complex tasks.
7 Conclusion
We studied the task of learning to play Othello and extended it to a multi-modal setting by introducing VisOthello. Our experiments examined whether access to visual state information improves sample efficiency and accelerates learning, comparing VisOthello against text-only and vision-only baselines. To further assess the benefits of multi-modal grounding, we introduce a board rotation perturbation and conduct feature alignment analysis to evaluate whether the models learn more robust and aligned representations. Our findings suggest that grounding language models with visual input leads to more efficient and stable learning. Beyond Othello, our framework provides a controlled testbed for the analysis of grounded representations and has the potential to extend to other model architectures, tasks, and modalities (see Appendix A for further discussion).
Limitations
Moreover, we do not include comparisons with large-scale text-only language models, as these have been thoroughly investigated in prior work Yuan and Søgaard (2025). Given that pretraining on language alone does not necessarily enhance understanding of the structured reasoning inherent in Othello, scaling up to such models and benchmarking against them is not currently a priority. Instead, our use of lightweight language models offers a practical and efficient probe into how much language pretraining contributes to this domain.
Ethics Statement
Acknowledgments
We thank the reviewers and area chair for their suggestions to incorporate the world model, which helped clarify our contributions better. We are also grateful to the members of the CoAStaL group at the University of Copenhagen and the IRLab at the University of Amsterdam for their valuable feedback on the experiments. Xinyi Chen is funded by the project LESSEN (NWA.1389.20.183) of the research program NWA-ORC 2020/21, which is (partly) financed by the Dutch Research Council (NWO), as well as by travel support from ELIAS (GA No.101120237) through the ELLIS program. Serge Belongie and Jiaang Li are supported by the Pioneer Centre for AI, DNRF grant number P1. Maarten de Rijke was supported by the Dutch Research Council (NWO), under project numbers 024.004.022, NWA.1389.20.183, and KICH3.LTP.20.006, and the European Union under grant agreements No. 101070212 (FINDHR) and No. 101201510 (UNITE). Views and opinions expressed are those of the author(s) only and do not necessarily reflect those of their respective employers, funders and/or granting authorities.
Appendix A Potential Impacts
While Othello is a harmless testbed, our methods for aligning multimodal features could, in principle, be adapted to sensitive domains. Our findings on data efficiency may also lower the barrier to training multimodal agents in low-resource settings. We emphasize that our work is intended solely for research on interpretability and grounding.
Appendix B Model Design Motivation
We use a masked language model (MLM) rather than an autoregressive multi-modal large language model (MLLM) for two main reasons detailed below.
Autoregressive training paradigm is not well-suited for our task setup.
Othello involves dynamic visual changes, as discs flip after each move (Figure 1). Understanding the current board state requires access to the complete move history, as it cannot be inferred from a single image alone—even for human players. Thus, the input must consist of a sequence of move tokens paired with the corresponding sequence of board states. This token-aligned multimodal sequence deviates significantly from standard MLLM training paradigms, which are typically designed for single image–text pairs or interleaved inputs without sequential dependencies. A more suitable framing is to model the game as a video sequence, with each board as a frame. However, feeding full sequences into current MLLMs introduces the risk of information leakage from future states and would require specialized causal multimodal masking, implying non-trivial architectural and training modifications.
Autoregressive training incurs extremely high resource costs.
Even if the technical challenges above were addressed, autoregressive training would remain computationally demanding: modeling an $n$-step game requires $n$ forward passes with progressively longer input sequences, whereas our MLM objective learns from the entire game in a single pass with partial masking. Our experiments were conducted on a single A100 GPU (40GB) with a dataset of approximately 20k Othello games, a scale that makes MLLM training infeasible.
Since the goal of this paper is to investigate the role of images in model understanding, we adopt an MLM-based approach with VisualBERT rather than MLLMs. This choice provides a lightweight framework for probing and analyzing the representations learned in multimodal Othello training.
Appendix C Model Training Details
All models are trained for up to 1000 epochs, with validation performed every 10 epochs. We apply early stopping with a patience of 5 validation steps, and retain the checkpoint with the highest validation accuracy for final evaluation. Training is conducted on a single NVIDIA A100-40GB GPU. BERT and VisOthello are trained with a batch size of 128 and a learning rate of 1e-4, while ResNet is trained with a batch size of 512 using the same learning rate.
Appendix D Model Sizes And Compute Resources
We report the parameter sizes and compute resources for all models used in our experiments. VisOthello, based on VisualBERT-base, has about 112M parameters. BERT (25 layers, hidden size 768) has about 177M parameters. Othello-GPT, based on GPT-2 Medium, contains 345M parameters. ResNet-18 has about 11.7M parameters. All models were trained on a single NVIDIA A100-40GB GPU. Training on 20k tasks required approximately 10 hours for VisOthello, 0.5 hours for ResNet-18, and 3.5 hours for BERT